OpenStudy (anonymous):

Arithmetic Sequence : If Sn = -441 for the series 19 + 15 + 11 + ... + tn, find n.

OpenStudy (anonymous):

sn = n/2(2a + (n-1)b) plug in the values of sn=-441, a=19, b=15-19=-4 we get : -441 = n/2 (2(19) + (n-1)*(-4)) solve for n !